# Phase Transitions in the Fluctuations of Functionals of Random Neural Networks

### 21 Apr 2026 v1

@@ 4.5 Non-central limit theorem
Let $\kappa$ satisfy 2 and set $\rho:=(\gamma_{1}-1)^{-1}$. There exist constants $\beta_{0}>0$ and $\beta_{1}$, depending only on $\kappa$, such that the renormalized iterates
@@
With the same assumptions and notations of Lemma 4.19, for fixed $q\in\mathbb{N}$ with $q\geq 2$, the following asymptotics hold.
@@
Case $\bm{q\geq\frac{d}{2\nu}}$. Let $C$ and $\delta$ are as in Lemma 4.19. Following the proof in the high-disorder case, we decompose the integral as
@@
The proof of the following corollary, which establishes Proposition 4.2 in the sparse case, partially follows the proof of the analogous result in the high-disorder case. However, the argument is more technical, since $1-C_{0}(L)$ converges to 0 rather than to a constant.
@@
Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume 2. Then for $q\geq 2$ we have the following.
@@
If $\beta\in L^{q}(\mu)$, then $\beta\in L^{k}(\mu)$ also for $k<q$ and so, $f_{L}\sim(\beta_{0}L)^{-1}$. The claim follows using Proposition 4.20. To prove that $\int_{-1}^{1}\widehat{\beta}^{q}(u)\,\mathrm{d}\mu(u)$ is always strictly positive, one can mirror exactly the argument in the proof of Corollary 4.15 to show that $\int_{-1}^{1}\widehat{\beta}^{q}(u)\,\mathrm{d}\mu(u)$ is 0 only if $\kappa$ is odd, just using Lemma 4.19 and (4.35) instead of Lemma 4.11 and (4.24). We conclude observing that $\kappa$ can never be odd in the sparse case (otherwise 0 would be a fixed point).
If $\beta\not\in L^{q}(\mu)$, i.e., $q\geq d/(2\nu)$, one can use the same argument as in the proof of Corollary 4.15 to prove that, for $k=1,\dots,q-1$, we have
@@
Then, the dominating term in (4.25) comes from $k=q$ and hence the claim. The proof of the upper bound when we consider $\int|\widehat{\kappa}_{L}|$ follows by standard inequalities and  Proposition 4.20. ∎
@@
Now, using the same computation as in Corollary 4.22, we have
@@ 5.1 Proof of Properties (P0)(P3)
5.2 Proof of Lemma 4.11: fixed-point analysis in the high-disorder regime
@@
From 1 we obtain $\kappa^{\prime}(1)-\kappa^{\prime}(x)\leq c_{\alpha}(1-x)^{\alpha}$ for $\alpha=\gamma-1>0$ and some $c_{\alpha}>0$. Thus, using the definition of $\nu$, we have $\kappa^{\prime}(1)=\lambda^{-1/\nu}$ and
@@
5.3 Proof of Lemma 4.4: bound for contractions
@@
5.4 Proof of Lemma 4.7: tail control in non-critical CLT
@@
Let 2 hold. Then for some $\delta\in(0,1)$ and $L$ large enough
@@
Using 2, there exists constants $\beth>0,\delta\in(0,1)$ such that for $x\in(0,\delta)$
@@
We prove that Lemma 4.7 holds both in the sparse and in the high-disorder regimes.
Proof of Lemma 4.7.
We observe that by arguing as in [38, Proof of Lemma 3.1], and using Proposition 4.2, we obtain
@@
where the last inequality follows from Proposition 4.2. ∎
5.5 Proof of Lemma 4.9: justification of the exchange of sum and integral
@@
5.6 Proof of Lemma 4.19: fixed-point analysis in the sparse regime.
To prove Lemma 4.19 we need the following auxiliary lemmas. We will sometimes use the notation
@@
In Lemma 4.19 we need a quantitative version of this convergence.
@@
Let $\kappa$ be not even and satisfy 2. Setting $A_{0}=c_{1}(\gamma_{1}-1)$ and $A_{1}=\frac{\gamma_{1}\,c_{1}}{2}+\frac{c_{2}}{c_{1}}\mathbf{1}_{\gamma_{2}=2\gamma_{1}-1}$ [^12] In this scenario, no assumption is made on the sign of $A_{1}$. For instance, if $\kappa$ is the kernel associated with the Gaussian activation function in the sparse case (see [20]), then $A_{1}=0$.. Then the renormalized iterates
@@
The analysis when $\kappa$ is even is analogous. Indeed the proof of Lemma 5.2 works also for $\kappa$ even, if one restrict $x\in[0,1)$. Since also $S_{L}$ is even we simply get
@@
First of all, by 2, there exist $D,\delta_{1}>0$ such that

### 23 Apr 2026 v2

@@ 4.5 Non-central limit theorem
Let $\kappa$ satisfy Assumption 2 and set $\rho:=(\gamma_{1}-1)^{-1}$. There exist constants $\beta_{0}>0$ and $\beta_{1}$, depending only on $\kappa$, such that the renormalized iterates
@@
With the same assumptions and notations of Section 4.7, for fixed $q\in\mathbb{N}$ with $q\geq 2$, the following asymptotics hold.
@@
Case $\bm{q\geq\frac{d}{2\nu}}$. Let $C$ and $\delta$ are as in Section 4.7. Following the proof in the high-disorder case, we decompose the integral as
@@
The proof of the following corollary, which establishes Section 4.2 in the sparse case, partially follows the proof of the analogous result in the high-disorder case. However, the argument is more technical, since $1-C_{0}(L)$ converges to 0 rather than to a constant.
@@
Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume Assumption 2. Then for $q\geq 2$ we have the following.
@@
If $\beta\in L^{q}(\mu)$, then $\beta\in L^{k}(\mu)$ also for $k<q$ and so, $f_{L}\sim(\beta_{0}L)^{-1}$. The claim follows using Section 4.7. To prove that $\int_{-1}^{1}\widehat{\beta}^{q}(u)\,\mathrm{d}\mu(u)$ is always strictly positive, one can mirror exactly the argument in the proof of Section 4.6 to show that $\int_{-1}^{1}\widehat{\beta}^{q}(u)\,\mathrm{d}\mu(u)$ is 0 only if $\kappa$ is odd, just using Section 4.7 and (4.35) instead of Section 4.6 and (4.24). We conclude observing that $\kappa$ can never be odd in the sparse case (otherwise 0 would be a fixed point).
If $\beta\not\in L^{q}(\mu)$, i.e., $q\geq d/(2\nu)$, one can use the same argument as in the proof of Section 4.6 to prove that, for $k=1,\dots,q-1$, we have
@@
Then, the dominating term in (4.25) comes from $k=q$ and hence the claim. The proof of the upper bound when we consider $\int|\widehat{\kappa}_{L}|$ follows by standard inequalities and  Section 4.7. ∎
@@
Now, using the same computation as in Section 4.7, we have
@@ 5.1 Proof of Properties (P0)(P3)
5.2 Proof of Section 4.6: fixed-point analysis in the high-disorder regime
@@
From Assumption 1 we obtain $\kappa^{\prime}(1)-\kappa^{\prime}(x)\leq c_{\alpha}(1-x)^{\alpha}$ for $\alpha=\gamma-1>0$ and some $c_{\alpha}>0$. Thus, using the definition of $\nu$, we have $\kappa^{\prime}(1)=\lambda^{-1/\nu}$ and
@@
5.3 Proof of Section 4.3: bound for contractions
@@
5.4 Proof of Section 4.4: tail control in non-critical CLT
@@
Let Assumption 2 hold. Then for some $\delta\in(0,1)$ and $L$ large enough
@@
Using Assumption 2, there exists constants $\beth>0,\delta\in(0,1)$ such that for $x\in(0,\delta)$
@@
We prove that Section 4.4 holds both in the sparse and in the high-disorder regimes.
Proof of Section 4.4.
We observe that by arguing as in [38, Proof of Lemma 3.1], and using Section 4.2, we obtain
@@
where the last inequality follows from Section 4.2. ∎
5.5 Proof of Section 4.5: justification of the exchange of sum and integral
@@
5.6 Proof of Section 4.7: fixed-point analysis in the sparse regime.
To prove Section 4.7 we need the following auxiliary lemmas. We will sometimes use the notation
@@
In Section 4.7 we need a quantitative version of this convergence.
@@
Let $\kappa$ be not even and satisfy Assumption 2. Setting $A_{0}=c_{1}(\gamma_{1}-1)$ and $A_{1}=\frac{\gamma_{1}\,c_{1}}{2}+\frac{c_{2}}{c_{1}}\mathbf{1}_{\gamma_{2}=2\gamma_{1}-1}$ [^12] In this scenario, no assumption is made on the sign of $A_{1}$. For instance, if $\kappa$ is the kernel associated with the Gaussian activation function in the sparse case (see [20]), then $A_{1}=0$.. Then the renormalized iterates
@@
The analysis when $\kappa$ is even is analogous. Indeed the proof of Section 5.6 works also for $\kappa$ even, if one restrict $x\in[0,1)$. Since also $S_{L}$ is even we simply get
@@
First of all, by Assumption 2, there exist $D,\delta_{1}>0$ such that